Memofax, Inc. produces memory enhancement software for computers. Sales have been very erratic, with some months showing a profit and some months showing a loss. The company’s contribution format income statement for the most recent month is given below:

  

  Sales (12,500 units at $20 per unit) $ 250,000   
  Less: Variable expenses 150,000   
  Contribution margin 100,000   
  Less: Fixed expenses 106,000   
  Net operating loss $ (6,000)
Required:
1.

Compute the company’s CM ratio and its break-even point in both units and dollars.

      

2.

The sales manager feels that an $10,000 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in a $80,000 increase in monthly sales. If the sales manager is right, what will be the effect on the company’s monthly net operating income or loss? (Use the incremental approach in preparing your answer.)

    

3.

Refer to the original data. The president is convinced that a 10% reduction in the selling price, combined with an increase of $46,000 in the monthly advertising budget, will double unit sales. What will the new contribution format income statement look like if these changes are adopted?

    

4.

Refer to the original data. The company’s advertising agency thinks that a new package would help sales. The new package being proposed would increase packaging costs by $0.5 per unit. Assuming no other changes, how many units would have to be sold each month to earn a profit of $6,500? (Do not round intermediate calculations.)

      

  

5.

Refer to the original data. By automating, the company could slash its variable expenses in half. However, fixed costs would increase by $92,800 per month.

a.

Compute the new CM ratio and the new break-even point in both units and dollars. (Do not round intermediate calculations. Round "Contribution Margin Ratio" to 2 decimal places.)

            

b.

Assume that the company expects to sell 20,000 units next month. Prepare two contribution format income statements: one assuming that operations are not automated, and one assuming that they are. (Do not round intermediate calculations. Round "Per Unit" and "Percentage" to 2 decimal places.)
Answer #1
contribution margin per unit= 100,000/12,500
8
1) CM ratio = contribution/sales
100,000/250,000
40.00%
BEP(units) = total fixed cost/contribution margin per unit
106000/8
13250
BEP(dollars) = 13250*20
265000
CM ratio 40%
Break even point in units 13250
Break even point in dollars 265000
2) increase in contribution (80,000*40%) 32000
less : increase in advertising budget 10,000
increase in net income 22,000
increases by 22,000
3) units = 12500*2 = 25000 units ; selling price = 20*90%=$18
Contribution Income statement
Sales (25000*18) 450000
Variable expense (25000*12) 300000
Contribution margin 150000
Fixed expenses (106000+46000) 152,000
Net income -2,000
4) New contribution margin = 8-.5
7.5
BEP(units) = (total fixed cost+target profit)/contribution per unit
(106000+6500)/7.5
15000
Sales units 15,000
5)
CM ratio = contribution/sales
14/20
70.00%
BEP(units) = total fixed cost/contribution margin per unit
(106000+92800)/14
14200
BEP(dollars) = 198800/70%
284000
CM ratio 70%
Break even point in units 14200
Break even point in dollars 284000
20000
b) Not Automated Automated
total per unit % total per unit %
Sales 400000 20 100% 400000 20 100%
Variable expenses 240000 12 60% 120000 6 30%
Contribution margin 160000 8 40% 280000 14 70%
Fixed expenses 106,000 198,800
Net operating income 54,000 81,200
c) yes